OpenStudy (anonymous):

1. Calculate the energy change (q) of the surroundings (water). We can assume that the specific heat capacity of water is 4.18 J / (g × °C) and the density of water is 1.00 g/mL. (27.776 g) * ( 4.18 J/g *C ) * ( 25.3 C ) Q=2944kJ ^^ is this correct? 2. Calculate the specific heat of the metal.

OpenStudy (aaronq):

use the formula: q=m*C*dT where dT=T(final)-T(initial)

OpenStudy (anonymous):

(27.776 g) * ( 4.18 J/g *C ) * ( 25.3 C ) Q=2944kJ Is this correct

OpenStudy (abb0t):

If you plug it in your calculator and that is your answer, then you should be correct.

OpenStudy (aaronq):

you were given the change in temperature, dT, as 25.3?

OpenStudy (anonymous):

Honestly I just need someone to walk me through this

OpenStudy (anonymous):

Measured Mass of Metal: 27.776g Distilled Water Volume: 26 mL Distilled Water Temperature: 25.3 *C Temperature of Metal: 100.5*C Temperature of Mixture:38.9*C 1. Calculate the energy change (q) of the surroundings (water). We can assume that the specific heat capacity of water is 4.18 J / (g × °C) and the density of water is 1.00 g/mL. qwater = m × c × ΔT m = mass of water = density x volume = 1 x 26 = 26 grams ΔT = T(mix) - T(water) = 38.9 - 25.3 = 13.6  q(water) = 26 x 13.6 x 4.18  q(water) = 1478 Joules 2. Calculate the specific heat of the metal. 3. q(water) = - q(metal) q(metal) = - 1478 Joules q(metal) = m × c × ΔT m = 27.776 g ΔT = T(mix) - T(metal)  ΔT = 38.9 - 100.5 = - 61.6  C = q(metal) / m x ΔT C = -1478 / (-61.6 x 27.776 ) C = 0.864 J / (g × °C)

OpenStudy (anonymous):

idk if I'm doing this correctly

OpenStudy (aaronq):

its good (Y)